# Yaroslav Aksenov, Ian Maksimov, Nikita Balagansky, Daniil Gavrilov Tinkoff

### v2

@@ D.3 Impact of Update Strategies on Gradient Dynamics in TR-DPO
Figure 14: The dynamics of gradient scales during TR-DPO training illustrate the influence of update strategies. Figure (a) shows gradient scale dynamics for TR-DPOα at $\beta=0.05$, reflecting the effect of soft updates with $\alpha=0.3,0.6,0.9$. Figure (b) illustrates the gradient scale dynamics for TR-DPOτ at $\beta=0.05$, detailing the gradient dynamics under hard updates with $\tau=64,256,512$. In both cases, the TR-DPO gradients are compared against the baseline DPO. See Section 4.4 for details.
Appendix E Diversity Analysis
Figures 15(a), 15(b), and 15(c) illustrate the relationship between mean HC metrics and Self-BLEU scores (Zhu et al.,, 2018) for the DPO, IPO, and KTO methods, respectively, adjusted by different $\beta$ values using the Pythia 2.8B model on the Anthropic-HH dataset. These graphs demonstrate that a decrease in $\beta$ leads to less diverse generations across all methods. Reinforcing the observations made by Wang et al., (2023), it is evident that higher alignment within models correlates with reduced diversity in generated outputs. The analysis confirms that at equivalent levels of response diversity, the TR-modified methods enhance HC metric values.
Figure 15: Figures (a), (b), and (c) show the relationship between the mean HC metrics and Self-BLEU scores for the DPO, IPO, and KTO methods respectively. Figure (a) illustrates results for DPO and TR-DPO with $\alpha=0.6$ and TR-DPO with $\tau=512$. Figure (b) shows the analysis for IPO and TR-IPO under similar conditions, and Figure (c) for KTO and TR-KTO, highlighting the effects of $\beta$ parameter adjustments. This analysis underscores the trend that lower $\beta$ values correspond to less diverse generations, affirming that models with higher alignment produce less varied outputs, and that TR modifications generally achieve higher HC metrics at the same level of diversity. See Section 4.4 for details.
In addition to diversity analysis, it is important to address the issue of the DPO method generating overly long texts, as identified by Park et al., (2024). The TR-DPO modification introduced in this study consistently produces shorter texts. Figure 16 demonstrates how output length varies with KL divergence for DPO, TR-DPOα ($\alpha=0.6$), and TR-DPOτ ($\tau=512$), across different $\beta$ values. This visualization highlights that, generally, TR-DPO produces shorter outputs than DPO at comparable levels of KL divergence, except TR-DPO with $\alpha=0.6$ and $\beta=0.01$, where an anomaly occurs due to the generation of repeated words.
@@ Appendix E Diversity Analysis
Appendix F GPT-4 AutoSxS details
@@
Table 3: AutoSxS results for DPO and TR-DPO with Fisher’s exact test for the Anthropic-HH and Reddit TL;DR datasets with Pythia 2.8B model. The results that are statistically significant in both tasks are highlighted in bold text.
Appendix G Generation examples
@@
G.1 Example from the Anthropic-HH Dataset
@@
G.2 Example from the Reddit TL;DR Dataset
@@
Appendix H GPT-4 evaluation prompts
@@
H.1 Pairwise Anthropic-HH Prompt
@@
H.2 Pairwise Reddit TL;DR summarization Prompt
@@
H.3 Human-Centric Anthropic-HH Prompt

### v3

@@ Appendix H Impact of Update Strategies on Gradient Dynamics in TR-DPO
Figure 17: The dynamics of exponentially smoothed gradient scales during TR-DPO training illustrate the influence of update strategies. Figure (a) shows gradient scale dynamics for TR-DPOα at $\beta=0.05$, reflecting the effect of soft updates with $\alpha=0.3,0.6,0.9$. Figure (b) illustrates the gradient scale dynamics for TR-DPOτ at $\beta=0.05$, detailing the gradient dynamics under hard updates with $\tau=64,256,512$. With a hard update, the real value of the gradient scale becomes 0.5, as the argument of the sigmoid becomes zero. The graph does not reflect this because smoothing of values is used for better visibility.
Appendix I Diversity Analysis
Figures 18(a), 18(b), and 18(c) illustrate the relationship between mean HC metrics and Self-BLEU scores (Zhu et al., 2018) for the DPO, IPO, and KTO methods, respectively, adjusted by different $\beta$ values using the Pythia 2.8B model on the Anthropic-HH dataset. These graphs demonstrate that a decrease in $\beta$ leads to less diverse generations across all methods. Reinforcing the observations made by Wang et al. (2023), it is evident that higher alignment within models correlates with reduced diversity in generated outputs. The analysis confirms that at equivalent levels of response diversity, the TR-modified methods enhance HC metric values.
Figure 18: Figures (a), (b), and (c) show the relationship between the mean HC metrics and Self-BLEU scores for the DPO, IPO, and KTO methods, respectively. Figure (a) illustrates results for DPO and TR-DPO with $\alpha=0.6$ and TR-DPO with $\tau=512$. Figure (b) shows the analysis for IPO and TR-IPO under similar conditions, and Figure (c) for KTO and TR-KTO, highlighting the effects of $\beta$ parameter adjustments. This analysis underscores the trend that lower $\beta$ values correspond to less diverse generations, affirming that models with higher alignment produce less varied outputs, and that TR modifications generally achieve higher HC metrics at the same level of diversity. See Section 4.4 for details.
In addition to diversity analysis, it is important to address the issue of the DPO method generating overly long texts, as identified by Park et al. (2024). The TR-DPO modification introduced in this study consistently produces shorter texts. Figure 19 demonstrates how output length varies with KL divergence for DPO, TR-DPOα ($\alpha=0.6$), and TR-DPOτ ($\tau=512$), across different $\beta$ values. This visualization highlights that, generally, TR-DPO produces shorter outputs than DPO at comparable levels of KL divergence, except TR-DPO with $\alpha=0.6$ and $\beta=0.01$, where an anomaly occurs due to the generation of repeated words.
@@ Appendix I Diversity Analysis
Appendix J GPT-4 AutoSxS Details
@@
Table 8: AutoSxS results for DPO and TR-DPO with Fisher’s exact test for the Anthropic-HH and Reddit TL;DR datasets with the Pythia 2.8B model. Statistically significant results in both tasks are highlighted in bold text.
Appendix K Generation Examples
@@
K.1 Example from the Anthropic-HH Dataset
@@
K.2 Example from the Reddit TL;DR Dataset
@@
Appendix L GPT-4 Evaluation Prompts
@@
L.1 Pairwise Anthropic-HH Prompt
@@
L.2 Pairwise Reddit TL;DR summarization Prompt
@@
L.3 Human-Centric Anthropic-HH Prompt